Aaron built a frame for his chicken pen. The length of the pen measures 2.3 m and the width of the pen measures 1.4m. The chicken wire comes in a pack that measures 2 m. How many packs will Aaron need to buy?


1
Expert's answer
2021-10-18T16:10:07-0400

Perimeter of the pen: P=(2.3+1.4)2=7.4  m.P=(2.3+1.4)*2=7.4\;m.

To find the number of packs we should round up 7.42=4.\frac{7.4}{2}=4.

Aaron need to buy 4 packs.